OpenStudy (anonymous):

sec^4 theta -sec^2 theta is equal to

OpenStudy (anonymous):

cos^2 theta-cos^4 theta is it right?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

no they are not equal you can tell just by plugging in any angle

OpenStudy (dumbcow):

what are you trying to do? simplify "sec^4 -sec^2 " ?

OpenStudy (anonymous):

\[\sec^4(\theta) - \sec^2(\theta) \implies \sec^2(\theta)[\sec^2(\theta) -1 ] \implies \sec^2(\theta) \cdot \tan^2(\theta)\]

OpenStudy (anonymous):

Or:\[\frac{1}{\cos^2(\theta)} \cdot \frac{\sin^2(\theta)}{\cos^2(\theta)} \implies \frac{\sin^2(\theta)}{\cos^4(\theta)}\]